Suppose in a simple economy with no foreign sector, the marginal propensity to save (MPS) is 0.25. Planned investment spending has suddenly fallen, reducing AE and output to a level that is $500 million below full employment level of income(Y*).
a) If the government decided to try to get the economy back to full employment by cutting taxes, how big of a tax cut would be needed? [5 points]
b) Using production, income and expenditure chain [graph your answer] explain adjustment process to full employment equilibrium. [10 points]
1
Expert's answer
2020-05-07T09:00:05-0400

a) mpc=1-mps=1-0.25=0.75

"mult=\frac{1}{0.25}=4"

"∆Y = -mpc \times(mult) \times ∆T"

"500 million = -0.75\times 4\times∆T"

"∆T = \frac{500}{-3}" million

 ∆T = -166.67 

(a tax cut of 166.67 million would be needed)

b)



tax cuts – growth in savings – investment growth – production growth – unemployment reduction –full employment equilibrium